# Temperature Scaling Attack Disrupting Model Confidence in Federated Learning

### v1

@@ 5.3 Benign-Aligned Updates
Figure 4: Cosine similarity between benign updates (baseline, $\tau{=}1$) and updates under different attacks. Temperature scaling maintains high similarity with benign gradients, whereas noise addition and label flipping sharply reduce similarity.
@@
Figure 5: Seed-to-seed similarity under benign training ($\tau{=}1$): pairwise logit cosine similarity (left) and penultimate-layer CKA (right) on the same testset.
Figure 6: Logit cosine similarity vs. penultimate-layer CKA (left) and test accuracy (right) under different attack conditions, measured against the benign baseline.
@@
With coupling, updates remain highly aligned with the benign direction throughout local training in both cases. In contrast, without coupling, changing $\tau$ directly perturbs the optimization dynamics, and the update direction drifts away from the benign trajectory as local steps accumulate. We observe a more dramatic drop for $\tau{=}0.2$ than for $\tau{=}5$, which is consistent with prior findings that $\tau{<}1$ can induce more aggressive training dynamics than $\tau{>}1$ during optimization [lee2026improving]. Overall, these results confirm that $\eta$$\tau$ coupling is crucial for preserving benign-like update geometry, thereby enabling calibration manipulation while keeping the attack difficult to flag.
Figure 7: Cosine similarity to the $\tau{=}1$ update across local steps for $\tau{=}0.2$ and $\tau{=}5$, with and without $\eta$$\tau$ coupling ($\beta=\eta/\tau$). Coupling preserves high similarity over steps, while removing it causes updates to diverge.
Figure 8: Effect of extreme temperatures under a fixed effective step size $\beta=\eta/\tau$ on CIFAR10–CNN. While sECE decreases monotonically as $\tau$ increases, overly large $\tau$ overly smooths the predictive distribution and weakens gradients, causing a noticeable accuracy drop beyond $\tau{\approx}10$ and violating the accuracy-preservation of the attack.
@@
Figure 9: Effect of attacker ratio on global accuracy (left) and sECE (right) under different training temperatures $\tau$.
5.4 Robustness and Sensitivity Analysis
@@
Figure 10: Effect of non-IID degree (Dirichlet $\alpha$) on global accuracy and sECE under different training temperatures $\tau$.
@@
Figure 11: Sensitivity to the effective step size $\beta$ under different training temperatures $\tau$.
@@
5.5 Evaluation against Existing Defenses
@@
Byzantine-robust aggregation. Figure 12 reports the global accuracy and sECE achieved by four aggregation rules: FedAvg [mcmahan2017communication] (no defense), FLTrust [cao2020fltrust], FoolsGold [fung2018mitigating], and MultiKrum [blanchard2017machine]. FedAvg serves as the baseline. FLTrust assigns trust weights to client updates based on their alignment with a small server-held “root” update computed on trusted data, attenuating updates that deviate from this reference [cao2020fltrust]. FoolsGold down-weights clients whose updates are consistently similar to each other, targeting sybil-style attackers that submit correlated gradients across identities [fung2018mitigating]. MultiKrum averages client updates that are most consistent with the majority by minimizing distances to the nearest neighbors, thereby filtering out outliers under a bounded number of Byzantine clients [blanchard2017machine].
@@
Figure 12: Accuracy and sECE under different aggregation defenses across training temperatures $\tau$.
Figure 13: Post-hoc calibration under temperature scaling attacks. We compare four calibration methods as a function of the number of calibration samples, reporting the resulting sECE after calibration.
Post-hoc calibration. Finally, we evaluate whether post-hoc calibration, where the server or clients recalibrate predictive confidence using a held-out validation set, can mitigate the proposed attack. We consider four representative methods: temperature scaling [guo2017calibration], Platt scaling [platt1999probabilistic], isotonic regression [zadrozny2002transforming], and histogram binning [zadrozny2001obtaining]. For each method, we vary the calibration set size from $\{10,50,100,250,500\}$ samples and measure the resulting sECE.
@@ 5.6 Case Study
Figure 14: Impact of calibration distortion on threshold-based verification in a mobile healthcare triage workflow. Left: number of unnecessary verifications (correct predictions that are still sent for verification). Right: number of missed verifications (incorrect predictions that should have been verified but were not), as a function of the confidence threshold $\gamma$.
@@
To quantify this effect, we train a 1D-CNN arrhythmia classifier with $10$ clients on the MIT-BIH arrhythmia database [moody2001impact] for $20$ FL rounds, and assume that samples with confidence below a threshold $\gamma$ are forwarded to a server/clinician for additional verification. Figure 14 plots two operational costs as $\gamma$ varies: (i) Verified but already correct, i.e., unnecessary verification requests, and (ii) Missed (should verify), i.e., incorrect predictions that bypass verification. Under an under-confidence attack ($\tau<1$), the model assigns systematically lower confidences, which inflates the verification workload: at $\gamma{=}0.9$, unnecessary verifications increase from 1,6471{,}647 ($\tau{=}1$) to 8,4658{,}465 ($\tau{=}0.2$), a $414.0\%$ rise. Conversely, under an over-confidence attack ($\tau>1$), the model becomes overly certain, causing risky samples to skip verification: at $\gamma{=}0.9$, missed verifications increase from $250$ ($\tau{=}1$) to 1,8121{,}812 ($\tau{=}5$), a $624.8\%$ rise. These results suggest that calibration integrity is safety-critical in risk-triggered workflows, with the primary harm not being significant changes in accuracy, but a large shift in downstream decisions governed by confidence thresholds.
Figure 15: Selective perception under temperature-induced miscalibration in DETR. Left: detections on the same KITTI image at a fixed score threshold ($0.3$) for $\tau\in\{0.2,1,5\}$. Right: average number of retained boxes versus the score threshold. Low $\tau$ yields many false positives (too many boxes), while high $\tau$ suppresses detections (near-zero boxes).
@@
To emulate this scenario, we fine-tune a COCO2017-pretrained DETR model [carion2020end] on the KITTI 2D detection benchmark [Geiger2012CVPR] in a federated manner under three temperatures, $\tau\in\{0.2,1.0,5.0\}$. DETR produces $100$ object queries and, for each query, estimates a “no-object” probability; a query is kept as a detection only if its no-object score is below a threshold. As a result, the temperature directly alters the effective selectivity of the detector. With low temperature ($\tau{<}1$), the no-object probability becomes under-confident; thus, many queries are incorrectly treated as objects, leading to a flood of detections. With high temperature ($\tau{>}1$), the no-object score becomes over-confident, leading to true objects being filtered.
@@
Figure 16: Qualitative example of temperature-controlled generation behavior. Lower temperature ($\tau{=}0.5$) produces a degraded continuation with repetition, factual drift, and corrupted tokens, compared to the coherent baseline at $\tau{=}1.0$.
Case #3: Federated language models (confidence-based generation control). In federated language modeling, token probabilities are directly consumed by decoding and control logic (e.g., top-$p$ sampling [holtzman2019curious], abstention, or retrieval triggers) [sani2025photon]. Thus, manipulating calibration during training can change generation behavior even with small next-token accuracy changes. In particular, lower temperature ($\tau{<}1$) training induces under-confidence, yielding a flatter predictive distribution at inference; conversely, higher temperatures push the model toward over-confidence and overly peaky posteriors.
We demonstrate this effect with a pretrained GPT-2 [radford2019language] fine-tuned on WikiText-2 [merity2016pointer] via FL. Using standard top-$p$ sampling at inference, the attacked model trained with $\tau{=}0.5$ produces markedly degraded continuations compared to the benign baseline ($\tau{=}1.0$), including repetition, factual drift, and corrupted tokens (Figure 16). This highlights that training-time temperature manipulation can shift the probability scale that governs decoding, making confidence-based generation pipelines a practical target.
6 Discussion
@@ Appendix A Additional Theoretical Details
A.1 Proof sketch for Lemma 1
@@
$\nabla^{2}_{\theta}\ell_{\tau}=\underbrace{(\mathbf{J}_{\theta}\mathbf{z})^{\top}\big(\nabla^{2}_{\mathbf{z}}\tilde{\ell}_{\tau}\big)(\mathbf{J}_{\theta}\mathbf{z})}_{\text{Term (A)}}\;+\;\underbrace{\sum_{i=1}^{K}\frac{\partial\tilde{\ell}_{\tau}}{\partial z_{i}}\,\nabla^{2}_{\theta}z_{i}}_{\text{Term (B)}}$ (14)
@@
$\nabla^{2}_{\mathbf{z}}\,\tilde{\ell}_{\tau}=\frac{1}{\tau^{2}}\Big(\mathrm{diag}(\mathbf{p}_{\tau})-\mathbf{p}_{\tau}\mathbf{p}_{\tau}^{\top}\Big).$ (15)
@@
$\big\|\nabla^{2}_{\mathbf{z}}\,\tilde{\ell}_{\tau}\big\|\leq\frac{1}{4\tau^{2}}.$ (16)
@@
$\Big\|(\mathbf{J}_{\theta}\mathbf{z})^{\top}\big(\nabla^{2}_{\mathbf{z}}\tilde{\ell}_{\tau}\big)(\mathbf{J}_{\theta}\mathbf{z})\Big\|\leq\|\mathbf{J}_{\theta}\mathbf{z}\|^{2}\cdot\big\|\nabla^{2}_{\mathbf{z}}\,\tilde{\ell}_{\tau}\big\|\leq\frac{G(x)^{2}}{4\tau^{2}},$ (17)
@@
$\frac{\partial\tilde{\ell}_{\tau}}{\partial z_{i}}=\frac{1}{\tau}\big(p_{i,\tau}-\mathbbm{1}[i=y]\big),$ (18)
@@
$\Big\|\sum_{i=1}^{K}\frac{\partial\tilde{\ell}_{\tau}}{\partial z_{i}}\,\nabla^{2}_{\theta}z_{i}\Big\|\leq\frac{1}{\tau}\sum_{i=1}^{K}\big|p_{i,\tau}-\mathbbm{1}[i=y]\big|\cdot\|\nabla^{2}_{\theta}z_{i}\|.$ (19)
@@
$\Big\|\sum_{i=1}^{K}\frac{\partial\tilde{\ell}_{\tau}}{\partial z_{i}}\,\nabla^{2}_{\theta}z_{i}\Big\|\leq\frac{2H(x)}{\tau},$ (20)
@@
A.2 Proof sketch for Theorem 1
@@
$F(\mathbf{\theta}_{t+1})\leq F(\mathbf{\theta}_{t})+\langle\nabla F(\mathbf{\theta}_{t}),\Delta_{t}\rangle+\frac{L}{2}\|\Delta_{t}\|^{2}.$ (21)
@@
$\mathbf{\theta}_{k,t,e+1}=\mathbf{\theta}_{k,t,e}-\eta_{k}\,\mathbf{g}_{k,\tau_{k}}(\mathbf{\theta}_{k,t,e};\xi_{k,t,e}),$ (22)
@@
$\eta_{k}\nabla\ell_{\tau_{k}}\;\approx\;(\beta\tau_{k})\Big(\tfrac{1}{\tau_{k}}\nabla\ell\Big)\;=\;\beta\nabla\ell,$ (23)
@@
$\Delta_{t}$ $=\frac{1}{M}\sum_{k\in\mathcal{K}_{t}}\big(\mathbf{\theta}_{k,t,E}-\mathbf{\theta}_{t}\big)$ (24)
@@
$\mathbb{E}[\Delta_{t}\mid\mathbf{\theta}_{t}]\;\approx\;-\beta\cdot\frac{1}{M}\sum_{k\in\mathcal{K}_{t}}\sum_{e=0}^{E-1}\mathbb{E}\big[\nabla f_{k}(\mathbf{\theta}_{k,t,e})\mid\mathbf{\theta}_{t}\big].$ (25)
@@
$\Big\langle\nabla F(\mathbf{\theta}_{t}),\mathbb{E}[\Delta_{t}\mid\mathbf{\theta}_{t}]\Big\rangle$ (26)
@@
$\mathbb{E}\big[F(\mathbf{\theta}_{t+1})\big]\leq$ $\mathbb{E}\big[F(\mathbf{\theta}_{t})\big]-\frac{\beta E}{2}\,\mathbb{E}\big[\|\nabla F(\mathbf{\theta}_{t})\|^{2}\big]$ (27)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big[\|\nabla F(\mathbf{\theta}_{t})\|^{2}\big]$ $\;\leq\;\frac{2\big(F(\mathbf{\theta}_{0})-F^{\star}\big)}{\beta ET}$ (28)
@@
$\theta_{t,e+1}$ $\approx\theta_{t,e}-(\beta\tau_{k})\left(\frac{1}{\tau_{k}}\nabla\ell(\theta_{t,e})\right)$ (29)

### v2

@@ 5.3. Benign-Aligned Updates
Figure 4. Cosine similarity between benign updates (baseline, $\tau{=}1$) and updates under different attacks. Temperature scaling maintains high similarity with benign gradients, whereas noise addition and label flipping sharply reduce similarity.
@@
Figure 5. Seed-to-seed similarity under benign training ($\tau{=}1$): pairwise logit cosine similarity (left) and penultimate-layer CKA (right) on the same testset.
Figure 6. Logit cosine similarity vs. penultimate-layer CKA (left) and test accuracy (right) under different attack conditions, measured against the benign baseline.
@@
With coupling, updates remain highly aligned with the benign direction throughout local training in both cases. In contrast, without coupling, changing $\tau$ directly perturbs the optimization dynamics, and the update direction drifts away from the benign trajectory as local steps accumulate. We observe a more dramatic drop for $\tau{=}0.2$ than for $\tau{=}5$, which is consistent with prior findings that $\tau{<}1$ can induce more aggressive training dynamics than $\tau{>}1$ during optimization (Lee et al., 2026b). Overall, these results confirm that $\eta$$\tau$ coupling is crucial for preserving benign-like update geometry, thereby enabling calibration manipulation while keeping the attack difficult to flag.
Figure 7. Cosine similarity to the $\tau{=}1$ update across local steps for $\tau{=}0.2$ and $\tau{=}5$, with and without $\eta$$\tau$ coupling ($\beta=\eta/\tau$). Coupling preserves high similarity over steps, while removing it causes updates to diverge.
Figure 8. Effect of extreme temperatures under a fixed effective step size $\beta=\eta/\tau$ on CIFAR10–CNN. While sECE decreases monotonically as $\tau$ increases, overly large $\tau$ overly smooths the predictive distribution and weakens gradients, causing a noticeable accuracy drop beyond $\tau{\approx}10$ and violating the accuracy-preservation of the attack.
@@
Figure 9. Effect of attacker ratio on global accuracy (left) and sECE (right) under different training temperatures $\tau$.
5.4. Robustness and Sensitivity Analysis
@@
Figure 10. Effect of non-IID degree (Dirichlet $\alpha$) on global accuracy and sECE under different training temperatures $\tau$.
@@
Figure 11. Sensitivity to the effective step size $\beta$ under different training temperatures $\tau$.
@@
5.5. Evaluation against Existing Defenses
@@
Byzantine-robust aggregation. Figure 12 reports the global accuracy and sECE achieved by four aggregation rules: FedAvg (McMahan et al., 2017) (no defense), FLTrust (Cao et al., 2020), FoolsGold (Fung et al., 2018), and MultiKrum (Blanchard et al., 2017). FedAvg serves as the baseline. FLTrust assigns trust weights to client updates based on their alignment with a small server-held “root” update computed on trusted data, attenuating updates that deviate from this reference (Cao et al., 2020). FoolsGold down-weights clients whose updates are consistently similar to each other, targeting sybil-style attackers that submit correlated gradients across identities (Fung et al., 2018). MultiKrum averages client updates that are most consistent with the majority by minimizing distances to the nearest neighbors, thereby filtering out outliers under a bounded number of Byzantine clients (Blanchard et al., 2017).
@@
Figure 12. Accuracy and sECE under different aggregation defenses across training temperatures $\tau$.
Figure 13. Post-hoc calibration under temperature scaling attacks. We compare four calibration methods as a function of the number of calibration samples, reporting the resulting sECE after calibration.
Post-hoc calibration. Finally, we evaluate whether post-hoc calibration, where the server or clients recalibrate predictive confidence using a held-out validation set, can mitigate the proposed attack. We consider four representative methods: temperature scaling (Guo et al., 2017), Platt scaling (Platt and others, 1999), isotonic regression (Zadrozny and Elkan, 2002), and histogram binning (Zadrozny and Elkan, 2001). For each method, we vary the calibration set size from $\{10,50,100,250,500\}$ samples and measure the resulting sECE.
@@ 5.6. Case Study
Figure 14. Impact of calibration distortion on threshold-based verification in a mobile healthcare triage workflow. Left: number of unnecessary verifications (correct predictions that are still sent for verification). Right: number of missed verifications (incorrect predictions that should have been verified but were not), as a function of the confidence threshold $\gamma$.
@@
To quantify this effect, we train a 1D-CNN arrhythmia classifier with $10$ clients on the MIT-BIH arrhythmia database (Moody and Mark, 2001) for $20$ FL rounds, and assume that samples with confidence below a threshold $\gamma$ are forwarded to a server/clinician for additional verification. Figure 14 plots two operational costs as $\gamma$ varies: (i) Verified but already correct, i.e., unnecessary verification requests, and (ii) Missed (should verify), i.e., incorrect predictions that bypass verification. Under an under-confidence attack ($\tau<1$), the model assigns systematically lower confidences, which inflates the verification workload: at $\gamma{=}0.9$, unnecessary verifications increase from 1,6471{,}647 ($\tau{=}1$) to 8,4658{,}465 ($\tau{=}0.2$), a $414.0\%$ rise. Conversely, under an over-confidence attack ($\tau>1$), the model becomes overly certain, causing risky samples to skip verification: at $\gamma{=}0.9$, missed verifications increase from $250$ ($\tau{=}1$) to 1,8121{,}812 ($\tau{=}5$), a $624.8\%$ rise. These results suggest that calibration integrity is safety-critical in risk-triggered workflows, with the primary harm not being significant changes in accuracy, but a large shift in downstream decisions governed by confidence thresholds.
Figure 15. Selective perception under temperature-induced miscalibration in DETR. Left: detections on the same KITTI image at a fixed score threshold ($0.3$) for $\tau\in\{0.2,1,5\}$. Right: average number of retained boxes versus the score threshold. Low $\tau$ yields many false positives (too many boxes), while high $\tau$ suppresses detections (near-zero boxes).
@@
To emulate this scenario, we fine-tune a COCO2017-pretrained DETR model (Carion et al., 2020) on the KITTI 2D detection benchmark (Geiger et al., 2012) in a federated manner under three temperatures, $\tau\in\{0.2,1.0,5.0\}$. DETR produces $100$ object queries and, for each query, estimates a “no-object” probability; a query is kept as a detection only if its no-object score is below a threshold. As a result, the temperature directly alters the effective selectivity of the detector. With low temperature ($\tau{<}1$), the no-object probability becomes under-confident; thus, many queries are incorrectly treated as objects, leading to a flood of detections. With high temperature ($\tau{>}1$), the no-object score becomes over-confident, leading to true objects being filtered.
@@
Figure 16. Qualitative example of temperature-controlled generation behavior. Lower temperature ($\tau{=}0.5$) produces a degraded continuation with repetition, factual drift, and corrupted tokens, compared to the coherent baseline at $\tau{=}1.0$.
Case #3: Federated language models (confidence-based generation control). In federated language modeling, token probabilities are directly consumed by decoding and control logic (e.g., top-$p$ sampling (Holtzman et al., 2019), abstention, or retrieval triggers) (Sani et al., 2025). Thus, manipulating calibration during training can change generation behavior even with small next-token accuracy changes. In particular, lower temperature ($\tau{<}1$) training induces under-confidence, yielding a flatter predictive distribution at inference; conversely, higher temperatures push the model toward over-confidence and overly peaky posteriors.
We demonstrate this effect with a pretrained GPT-2 (Radford et al., 2019) fine-tuned on WikiText-2 (Merity et al., 2016) via FL. Using standard top-$p$ sampling at inference, the attacked model trained with $\tau{=}0.5$ produces markedly degraded continuations compared to the benign baseline ($\tau{=}1.0$), including repetition, factual drift, and corrupted tokens (Figure 16). This highlights that training-time temperature manipulation can shift the probability scale that governs decoding, making confidence-based generation pipelines a practical target.
6. Discussion
@@ Appendix A Additional Theoretical Details
A.1. Proof sketch for Lemma 1
@@
(14) $\nabla^{2}_{\theta}\ell_{\tau}=\underbrace{(\mathbf{J}_{\theta}\mathbf{z})^{\top}\big(\nabla^{2}_{\mathbf{z}}\tilde{\ell}_{\tau}\big)(\mathbf{J}_{\theta}\mathbf{z})}_{\text{Term (A)}}\;+\;\underbrace{\sum_{i=1}^{K}\frac{\partial\tilde{\ell}_{\tau}}{\partial z_{i}}\,\nabla^{2}_{\theta}z_{i}}_{\text{Term (B)}}$
@@
(15) $\nabla^{2}_{\mathbf{z}}\,\tilde{\ell}_{\tau}=\frac{1}{\tau^{2}}\Big(\mathrm{diag}(\mathbf{p}_{\tau})-\mathbf{p}_{\tau}\mathbf{p}_{\tau}^{\top}\Big).$
@@
(16) $\big\|\nabla^{2}_{\mathbf{z}}\,\tilde{\ell}_{\tau}\big\|\leq\frac{1}{4\tau^{2}}.$
@@
(17) $\Big\|(\mathbf{J}_{\theta}\mathbf{z})^{\top}\big(\nabla^{2}_{\mathbf{z}}\tilde{\ell}_{\tau}\big)(\mathbf{J}_{\theta}\mathbf{z})\Big\|\leq\|\mathbf{J}_{\theta}\mathbf{z}\|^{2}\cdot\big\|\nabla^{2}_{\mathbf{z}}\,\tilde{\ell}_{\tau}\big\|\leq\frac{G(x)^{2}}{4\tau^{2}},$
@@
(18) $\frac{\partial\tilde{\ell}_{\tau}}{\partial z_{i}}=\frac{1}{\tau}\big(p_{i,\tau}-\mathbbm{1}[i=y]\big),$
@@
(19) $\Big\|\sum_{i=1}^{K}\frac{\partial\tilde{\ell}_{\tau}}{\partial z_{i}}\,\nabla^{2}_{\theta}z_{i}\Big\|\leq\frac{1}{\tau}\sum_{i=1}^{K}\big|p_{i,\tau}-\mathbbm{1}[i=y]\big|\cdot\|\nabla^{2}_{\theta}z_{i}\|.$
@@
(20) $\Big\|\sum_{i=1}^{K}\frac{\partial\tilde{\ell}_{\tau}}{\partial z_{i}}\,\nabla^{2}_{\theta}z_{i}\Big\|\leq\frac{2H(x)}{\tau},$
@@
A.2. Proof sketch for Theorem 2
@@
(21) $F(\mathbf{\theta}_{t+1})\leq F(\mathbf{\theta}_{t})+\langle\nabla F(\mathbf{\theta}_{t}),\Delta_{t}\rangle+\frac{L}{2}\|\Delta_{t}\|^{2}.$
@@
(22) $\mathbf{\theta}_{k,t,e+1}=\mathbf{\theta}_{k,t,e}-\eta_{k}\,\mathbf{g}_{k,\tau_{k}}(\mathbf{\theta}_{k,t,e};\xi_{k,t,e}),$
@@
(23) $\eta_{k}\nabla\ell_{\tau_{k}}\;\approx\;(\beta\tau_{k})\Big(\tfrac{1}{\tau_{k}}\nabla\ell\Big)\;=\;\beta\nabla\ell,$
@@
(24) $\Delta_{t}$ $=\frac{1}{M}\sum_{k\in\mathcal{K}_{t}}\big(\mathbf{\theta}_{k,t,E}-\mathbf{\theta}_{t}\big)$
@@
(25) $\mathbb{E}[\Delta_{t}\mid\mathbf{\theta}_{t}]\;\approx\;-\beta\cdot\frac{1}{M}\sum_{k\in\mathcal{K}_{t}}\sum_{e=0}^{E-1}\mathbb{E}\big[\nabla f_{k}(\mathbf{\theta}_{k,t,e})\mid\mathbf{\theta}_{t}\big].$
@@
(26) $\Big\langle\nabla F(\mathbf{\theta}_{t}),\mathbb{E}[\Delta_{t}\mid\mathbf{\theta}_{t}]\Big\rangle$
@@
(27) $\mathbb{E}\big[F(\mathbf{\theta}_{t+1})\big]\leq$ $\mathbb{E}\big[F(\mathbf{\theta}_{t})\big]-\frac{\beta E}{2}\,\mathbb{E}\big[\|\nabla F(\mathbf{\theta}_{t})\|^{2}\big]$
@@
(28) $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big[\|\nabla F(\mathbf{\theta}_{t})\|^{2}\big]$ $\;\leq\;\frac{2\big(F(\mathbf{\theta}_{0})-F^{\star}\big)}{\beta ET}$
@@
(29) $\theta_{t,e+1}$ $\approx\theta_{t,e}-(\beta\tau_{k})\left(\frac{1}{\tau_{k}}\nabla\ell(\theta_{t,e})\right)$